OpenStudy (anonymous):

In a constant-pressure calorimeter, 55.0 mL of 0.330 M Ba(OH)2 was added to 55.0 mL of 0.660 M HCl. The reaction caused the temperature of the solution to rise from 24.23 °C to 28.73 °C. If the solution has the same density and specific heat as water, what is ΔH for this reaction (per mole of H2O produced)? Assume that the total volume is the sum of the individual volumes.

OpenStudy (jhannybean):

\[\sf Ba(OH)_2 ~(aq)~ + 2HCl~(aq) ~\rightarrow~ BaCl_2 ~(aq)~+~2H_2O~(l)\]

OpenStudy (jhannybean):

\[\sf 0.0550~L~\times~\frac{0.330~mol~Ba(OH)_2}{1~L} =0.0182~mol~Ba(OH)_2 \]\[\sf 0.0550~L~\times~\frac{0.660~mol~HCl}{1~L} = 0.0363~mol~HCl\] Guessing you use \(\sf q=mc\Delta t\) to find \(\sf q\), then to find \(\sf \Delta H_f\) you use \(\sum {n \cdot \text{prod} -\sum m\cdot \text{reactants}}\) ? My guess....